OpenStudy (anonymous):

Rewrite the number in scientific notation: 0.02

OpenStudy (anonymous):

move decimal point to the right two places, you get \(2\times 10^{-2}\)

OpenStudy (anonymous):

or you can think \(0.02=2(0.01)=2\times \frac{1}{100}=2\times 10^{-2}\)

OpenStudy (anonymous):

2 x 10^-2